OpenStudy (anonymous):

Solve: |x + 3| = 3 A. {0, 6} B. {0, -6} C. {-6, 6} D. {-6}

OpenStudy (hoblos):

|x + 3| = 3 x+3 = 3 OR x+3=-3 x=0 x=-6 {0, -6}

OpenStudy (hoblos):

|3r - 12| = 0 3r -12 = 0 3r=12 r=4 D